# Non-minimal light-curvature couplings and black-hole imaging

### v1

@@ Abstract
Non-minimal couplings between the electromagnetic field strength and the spacetime curvature are part of the effective field theory of gravity and matter. They alter the local propagation of light in a significant way if the ratio of spacetime curvature to the non-minimal coupling is of order one. Spacetime curvature can become appreciable around black holes, and yet the effect of non-minimal couplings on electromagnetic observations of black holes remains underexplored. A particular feature of the non-minimal coupling between the electromagnetic field-strength and the Riemann tensor is that it generates two distinct photon rings for different polarizations. Working within the paradigm of lensing bands, and focusing on the $n=1$ photon ring, we investigate which values of the non-minimal coupling can be robustly excluded.
@@ I Introduction
Non-minimal interactions between light and spacetime curvature are ubiquitous. For instance, they appear in the Euler-Heisenberg action on a curved background, i.e., they are generated in the low-energy regime of the Einstein-Maxwell theory by integrating out quantum fluctuations of the electron [1, 2]. Even at a purely classical level, non-minimal interactions are allowed as part of the most general Lagrangian for gravitational and electromagnetic fields that leads to second-order equations of motion [3, 4].
@@
The non-minimal interactions are built from various contractions of the electromagnetic field strength $F_{\mu\nu}$ with the Riemann tensor $R_{\mu\nu\kappa\lambda}$ and its own contractions, as well as with the Levi-Civita tensor $\epsilon_{\mu\nu\rho\sigma}$. Possible interactions include, e.g., $F_{\mu\nu}F_{\rho\sigma}C^{\mu\nu\rho\sigma}$ with the Weyl tensor $C^{\mu\nu\rho\sigma}$, which was recently investigated in quantum gravity [5]. Another example is the Horndeski term, which couples the double-dual of the Riemann tensor to two powers of the electromagnetic field strength [3, 4].
@@
The lowest-order non-minimal interactions are dimension-six operators and, in accordance with standard effective-field-theory (EFT) arguments, are expected to be suppressed by the square of the mass-scale of the particle that is integrated out, e.g., the electron. If a very light and milli-charged particle exists [^1] See [6, 7, 8, 9, 10] for up-to-date studies of the parameter space., then the corresponding mass scale can be much lower, rendering those operators less suppressed at lower energies. From a different perspective, positivity bounds constrain such couplings [11], such that the EFT does not allow arbitrary coefficients.
@@
It is also expected that quantum gravitational fluctuations generate such terms. Asymptotically safe quantum gravity is an explicit example of a candidate theory that gives rise to such terms [5]. Also within a quantum-gravitational context, the couplings between light and curvature are restricted by the weak-gravity conjecture [11, 5].
@@
From a more phenomenological perspective, non-minimal couplings could be relevant in settings in which light propagates on high-curvature backgrounds [1, 12, 13, 14, 15, 16]. Important examples are black-hole images, which arise when light propagates on a black-hole spacetime. Following (and in part even preceding) the first images of M87* [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] and Sgr A* [27, 28, 29, 30, 31, 32, 33, 34] and their subsequent observations [35, 36], numerous studies explore the potential to constrain deviations from General Relativity (GR), see, e.g., [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 32, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] and references therein. In these studies, it is typically assumed that the parameters that control the deviation from GR are large, i.e., the dimensionless ratio of beyond-GR couplings to an appropriate power of the gravitational radius is typically assumed to be $\mathcal{O}(1)$. From an EFT perspective and invoking a naturalness argument, such a setting should also feature non-minimal couplings. However, effects that arise from non-minimal couplings have been significantly less studied [79, 80, 16, 81]. In building a physics-case for future VLBI-arrays, either ground-based [82, 83, 71] or space-based [84, 85, 86], it is important to understand which new-physics effects can (i) be tested in principle and (ii) be distinguished from each other (see, e.g., [87, 88, 89, 90]). For this goal, it is important to figure out whether non-minimal couplings, which modify the propagation of light on a curved spacetime, can have similar effects to the modification of the spacetime itself. With this motivation, we explore the impact of a non-minimal photon-curvature coupling on black-hole image features.
@@ II Non-minimal light-curvature couplings and effective field theory
Non-minimal couplings of the electromagnetic field strength $F_{\mu\nu}$ to the spacetime curvature without negative powers of derivatives are necessarily at least dimension-six operators and contain four derivatives, because otherwise $U(1)$ gauge symmetry and/or diffeomorphism invariance cannot be respected. Models with more than two time derivatives are difficult to treat, because, according to the Ostrogradsky theorem from classical mechanics, such field theories have a Hamiltonian that is unbounded from below under a non-degeneracy assumption. [^2] It is usually assumed that such an unbounded Hamiltonian results in a catastrophic instability of the theory. Recently, this assumption has been challenged by explicit counterexamples as well as proofs of well-defined time-evolution [91, 92, 93, 94]. A full theory of when higher-order time derivatives do not lead to catastrophic instabilities has not yet been reached. Once it has become available, we expect that a larger subset of EFTs for photon-curvature interactions will become accessible. Thus, we focus on the unique non-minimal interaction that can be added to the Einstein-Maxwell Lagrangian without resulting in higher-order equations of motion. This interaction was derived by Horndeski [3] and leads to the Horndeski vector-tensor theory given by the Lagrangian [3, 4]
@@
is the double-dual Riemann tensor, $\epsilon_{\mu\nu\alpha\beta}$ the Levi-Civita tensor and $\alpha$ is a coupling constant with dimensions of length squared. [^3] We work in units in which $\hbar=1=c$. The Horndeski term can also be expressed as
@@
The Horndeski vector-tensor term also emerges in the effective four-dimensional theory obtained by a Kaluza-Klein reduction of higher-dimensional Gauss-Bonnet gravity [95].
@@
The above Lagrangian should not be understood as a proposal for a fundamental theory. It may rather be understood as an EFT-description of the gravity-vector system in which the coupling $\alpha$ is parametrically larger than the other couplings in the fourth order of the derivative expansion. [^4] Often in EFT treatments, it is assumed that dimensionless couplings are of $\mathcal{O}(1)$, such that all dimensionful couplings are set by the same mass scale. This need not in general be the case, because the microscopic theory may contain mechanisms to selectively enhance just a subset of the couplings in the EFT compared to the other couplings at the same order of the derivative expansion. Thus, the above Lagrangian may even be a complete EFT (to that order). However, it is best understood as a toy model, in which we can explore which types of effects may arise due to non-minimal interactions and we choose the model guided by technical simplicity, in particular requiring second-order equations of motion. As we will see, the model is already complex enough to exhibit several distinct effects in black-hole images.
@@
Previously, Ref. [96] derived bounds on the non-minimal coupling $\alpha$. From the birefringent bending of light by the sun, Ref. [96] obtained $\sqrt{|\alpha|}\lesssim 3.16\times 10^{5}\,\mathrm{km}$, while from radar echo experiments of radio signals passing in the vicinity of the sun they obtained $\sqrt{|\alpha|}\lesssim 6.25\times 10^{4}\,\mathrm{km}$, and $\sqrt{|\alpha|}\lesssim 2.45\,\mathrm{km}$ from the timing of the binary pulsar PSR B1534+12 signals. Taking these constraints at face value would rule out $|\alpha|/M^{2}\sim\mathcal{O}(1)$ (for $M$ the mass of a supermassive black hole), which is the range of values for which we expect black-hole images to change appreciably, seemingly making our study superfluous. This is, however, not the case, for the following two reasons. First, we caution that astrophysical constraints are typically subject to model-dependence, and the present constraints are no exception. For instance, in different astrophysical environments, both magnetic fields as well as a non-vanishing Ricci scalar are present. Both affect the propagation of light, if higher-order field-strength terms as well as a non-minimal Ricci coupling are included in the Lagrangian. Translating observations into constraints is subject to assumptions about whether or not such couplings are present in Lagrangians. Second, we consider $\alpha$, rather than other (so far unconstrained) non-minimal couplings, purely for reasons of calculational convenience, because $\alpha$, being part of the Horndeski family and having an effective metric description for photon propagation, c.f. Sec. III, allows us to compute effects easily. We expect that the effects induced by $\alpha$ are paradigmatic of those created by non-minimal couplings more generally, motivating our study of their imprints on black-hole images.
@@ III Propagation of non-minimally coupled photons
In the presence of a non-minimal coupling, the propagation of photons is in general not described by the standard geodesic equation [1]. The latter is obtained from the wave-like solutions of motion for minimally coupled electromagnetic fields in the geometric-optics limit. The equation of motion is modified by non-minimal interactions; therefore the null geodesics of the background spacetime $g_{\mu\nu}$ lose their interpretation as describing the propagation of photons.
At a technical level, some settings with non-minimal couplings enable a description through the geodesic equation of an effective metric $\tilde{g}_{\mu\nu}$. This effective metric does not encode the geometry of spacetime, instead, it captures the effect of the non-minimal interaction at an effective level. This can be done if there is a field transformation for the particular solution $g_{\mu\nu}\rightarrow\tilde{g}_{\mu\nu}$ that maps the Lagrangian in Eq. (1) into the Einstein-Maxwell Lagrangian plus higher-curvature corrections. However, as we will see below, the situation for black-hole backgrounds is more complex and, in particular, a single effective metric does not suffice.
@@
where the square brackets correspond to the anti-symmetrization of indices as $T_{[\mu\nu]}=(T_{\mu\nu}-T_{\nu\mu})/2$. Eq. (7) is a homogeneous system of equations for the tensor $f_{\mu\nu}$ that admits non-trivial solutions if and only if $k^{\mu}$ is suitably constrained. To extract the form of the relevant constraints, we note that the Riemann tensor for the Schwarzschild spacetime can be written as [1]
@@
where $\{U_{(i)}^{\mu\nu}\}_{i=1}^{2}$ are defined in terms of vierbein fields $\{e^{\mu}_{(a)}\}_{a=0}^{3}$ satisfying $g^{\mu\nu}=\eta^{ab}e^{\mu}_{(a)}e^{\mu}_{(b)}$, with $\eta^{ab}$ the inverse flat metric, as [1]
@@
It holds that $k^{\mu}l_{\mu}=0=k^{\mu}m_{\mu}$, so that $l_{\mu}$ and $m_{\mu}$ can be understood as the two basis vectors which encode the polarization of a photon in the plane orthogonal to its direction of propagation. The corresponding effective metrics can be written (up to an arbitrary conformal factor, because the effective metrics only encode modified propagation of light) as [79]
@@
$-1/2\leq\alpha/M^{2}\leq 1,$ (15)
in order for the effective metrics to be non-singular for both polarizations at the horizon $r=2M$.
In practice, the effective metrics above indicate that light rays do not propagate along null geodesics of the metric $g_{\mu\nu}$. Photons with different polarizations propagating on the Schwarzschild spacetime experience different effective metrics in the geometric optics approximation (later studies also considered the cases of charged and rotating black-hole spacetimes [12, 13] and obtained the effective metrics in the geometric optics approximation, see, e.g., [80]). This is a phenomenon known as bi-metricity [97]. In the sections below, we will determine the impact this feature can have in black-hole images.

### v2

@@ Abstract
Non-minimal couplings between the electromagnetic field strength and the spacetime curvature are part of the effective field theory of gravity and matter. They alter the local propagation of light in a significant way if the ratio of spacetime curvature to the non-minimal coupling is of order one. Spacetime curvature can become appreciable around black holes, and yet the effect of non-minimal couplings on electromagnetic observations of black holes remains underexplored. A particular feature of the non-minimal coupling between the electromagnetic field-strength and the Riemann tensor is that it generates two distinct photon rings for different polarizations. Working within the paradigm of lensing bands and focusing on the $n=1$ lensing band, we illustrate by which diagnostics a modified light propagation may be distinguished from a modified spacetime geometry and how constraints on the value of the non-minimal coupling can be obtained.
@@ I Introduction
Non-minimal interactions between light and spacetime curvature are ubiquitous. For instance, they appear in the Euler-Heisenberg action on a curved background, i.e., they are generated in the low-energy regime of the Einstein-Maxwell theory by integrating out quantum fluctuations of the electron [Drummond:1979pp, Bastianelli:2008cu]. Even at a purely classical level, non-minimal interactions are allowed as part of the most general Lagrangian for gravitational and electromagnetic fields that leads to second-order equations of motion [Horndeski:1976gi, BeltranJimenez:2013btb].
@@
The non-minimal interactions are built from various contractions of the electromagnetic field-strength $F_{\mu\nu}$ with the Riemann tensor $R_{\mu\nu\kappa\lambda}$ and its own contractions, as well as with the Levi-Civita tensor $\epsilon_{\mu\nu\rho\sigma}$. Possible interactions include, e.g., $F_{\mu\nu}F_{\rho\sigma}C^{\mu\nu\rho\sigma}$ with the Weyl tensor $C^{\mu\nu\rho\sigma}$, which was recently investigated in quantum gravity [Knorr:2024yiu]. Another example is the Horndeski term, which couples the double-dual of the Riemann tensor to two powers of the electromagnetic field-strength [Horndeski:1976gi, BeltranJimenez:2013btb].
@@
The lowest-order non-minimal interactions are dimension-six operators and, in accordance with standard effective-field-theory (EFT) arguments, are expected to be suppressed by the square of the mass-scale of the particle that is integrated out, e.g., the electron. If a very light and milli-charged particle exists [^1] See [Davidson:2000hf, Boyarsky:2018tvu, Alvey:2020xsk, deMontigny:2023qft, Kalliokoski:2023cgw] for up-to-date studies of the parameter space., then the corresponding mass scale can be much lower, rendering those operators less suppressed at lower energies. From a different perspective, positivity bounds constrain such couplings [Bellazzini:2019xts], such that the EFT does not allow arbitrary coefficients.
@@
It is also expected that quantum gravitational fluctuations generate such terms. Asymptotically safe quantum gravity is an explicit example of a candidate theory that gives rise to such terms [Knorr:2024yiu]. Also within a quantum-gravitational context, the couplings between light and curvature are restricted by the weak-gravity conjecture [Bellazzini:2019xts, Knorr:2024yiu].
@@
From a more phenomenological perspective, non-minimal couplings could be relevant in settings in which light propagates on high-curvature backgrounds [Drummond:1979pp, Daniels:1993yi, Daniels:1995yw, Shore:1995fz, Shore:2002gn, Jana:2021lqe]. Important examples are black-hole images, which arise when light propagates on a black-hole spacetime. Following (and in part even preceding) the first images of M87* [EventHorizonTelescope:2019dse, EventHorizonTelescope:2019uob, EventHorizonTelescope:2019jan, EventHorizonTelescope:2019ths, EventHorizonTelescope:2019pgp, EventHorizonTelescope:2019ggy, EventHorizonTelescope:2021bee, EventHorizonTelescope:2021dqv, EventHorizonTelescope:2021srq, EventHorizonTelescope:2023gtd] and Sgr A* [EventHorizonTelescope:2022wkp, EventHorizonTelescope:2022apq, EventHorizonTelescope:2022wok, EventHorizonTelescope:2022exc, EventHorizonTelescope:2022urf, EventHorizonTelescope:2022xqj, EventHorizonTelescope:2024hpu, EventHorizonTelescope:2024rju] and their subsequent observations [EventHorizonTelescope:2024dhe, EventHorizonTelescope:2024uoo], numerous studies explore the potential to constrain deviations from General Relativity (GR), see, e.g., [Broderick:2013rlq, Wei:2013kza, Grenzebach:2014fha, Johannsen:2015hib, Cunha:2015yba, Cunha:2016wzk, Vincent:2016sjq, Abdujabbarov:2016hnw, Amir:2016cen, Tsukamoto:2017fxq, Ayzenberg:2018jip, Psaltis:2018xkc, Mizuno:2018lxz, Cunha:2018acu, Held:2019xde, Tian:2019yhn, Vagnozzi:2019apd, Allahyari:2019jqz, Cunha:2019dwb, Volkel:2020xlc, Konoplya:2020bxa, Khodadi:2021gbc, Younsi:2021dxe, EventHorizonTelescope:2022xqj, Vagnozzi:2022moj, Eichhorn:2022fcl, Carballo-Rubio:2022aed, Chen:2022scf, Khodadi:2020jij, Fernandes:2024ztk, dePaula:2023ozi, Sengo:2022jif, Junior:2021dyw, Xavier:2020egv, Olmo:2023lil, Ayzenberg:2023hfw, Carballo-Rubio:2023ekp, Afrin:2022ztr, Salehi:2023eqy, Lupsasca:2024wkp, Broderick:2024vjp, Khodadi:2024ubi, Carballo-Rubio:2025fnc, Bambi:2025wjx] and references therein. In these studies, it is typically assumed that the parameters that control the deviation from GR are large, i.e., the dimensionless ratio of beyond-GR couplings to an appropriate power of the gravitational radius is typically assumed to be $\mathcal{O}(1)$. From an EFT perspective and invoking a naturalness argument, such a setting should also feature non-minimal couplings. However, effects that arise from non-minimal couplings have been significantly less studied [Chen:2015cpa, Chen:2016hil, Jana:2021lqe, Chen:2023wna]. In building a physics-case for future Very Long Baseline Interferometry (VLBI)-arrays, either ground-based [Doeleman:2023kzg, Johnson:2023ynn, Ayzenberg:2023hfw] or space-based [2024SPIE13092E..6PI, Johnson:2024ttr, Zineb:2024gwx], it is important to understand which new-physics effects can (i) be tested in principle and (ii) be distinguished from each other (see, e.g., [Glampedakis:2021oie, Lara:2021zth, Ozel:2021ayr, Kocherlakota:2022jnz]). For this goal, it is important to figure out whether non-minimal couplings, which modify the propagation of light on a curved spacetime, can have similar effects to the modification of the spacetime itself. With this motivation, we explore the impact of a non-minimal photon-curvature coupling on black-hole image features.
@@ II Non-minimal light-curvature couplings and effective field theory
Non-minimal couplings of the electromagnetic field-strength $F_{\mu\nu}$ to the spacetime curvature without negative powers of derivatives are necessarily at least dimension-six operators and contain four derivatives, because otherwise $U(1)$ gauge symmetry and/or diffeomorphism invariance cannot be respected. Models with more than two time derivatives are difficult to treat, because, according to the Ostrogradsky theorem from classical mechanics, such field theories have a Hamiltonian that is unbounded from below under a non-degeneracy assumption. [^2] It is usually assumed that such an unbounded Hamiltonian results in a catastrophic instability of the theory. Recently, this assumption has been challenged by explicit counterexamples as well as proofs of well-defined time-evolution [Deffayet:2021nnt, Deffayet:2023wdg, Deffayet:2025lnj, Held:2025ckb]. A full theory of when higher-order time derivatives do not lead to catastrophic instabilities has not yet been reached. Once it has become available, we expect that a larger subset of EFTs for photon-curvature interactions will become accessible. Thus, we focus on the unique non-minimal interaction that can be added to the Einstein-Maxwell Lagrangian without resulting in higher-order equations of motion. This interaction was derived by Horndeski [Horndeski:1976gi] and leads to the Horndeski vector-tensor theory given by the Lagrangian [Horndeski:1976gi, BeltranJimenez:2013btb]
@@
is the double-dual Riemann tensor, $\epsilon_{\mu\nu\alpha\beta}$ the Levi-Civita tensor and $\alpha$ is a coupling constant with dimensions of length squared [^3] We work in units in which $\hbar=1=c$. controlling the strength of the non-minimal coupling to gravity. [^4] Although the strength and sign of $\alpha$ are arbitrary, they might be constrained by stability requirements. For an exterior Schwarzschild background to be stable, $\alpha$ can take any sign, but must be subdominant compared to the minimal coupling term, see [BeltranJimenez:2013btb]. The Horndeski term can also be expressed as
@@
The Horndeski vector-tensor term also emerges in the effective four-dimensional theory obtained by a Kaluza-Klein reduction of higher-dimensional Gauss-Bonnet gravity [Buchdahl:1979wi].
@@
The above Lagrangian should not be understood as a proposal for a fundamental theory. It may rather be understood as an EFT-description of the gravity-vector system in which the coupling $\alpha$ is parametrically larger than the other couplings in the fourth order of the derivative expansion. [^5] Often in EFT treatments, it is assumed that dimensionless couplings are of $\mathcal{O}(1)$, such that all dimensionful couplings are set by the same mass scale. This need not in general be the case, because the microscopic theory may contain mechanisms to selectively enhance just a subset of the couplings in the EFT compared to the other couplings at the same order of the derivative expansion. One example of such a mechanism relies on non-trivial scaling dimensions which arise due to quantum fluctuations, see, e.g., [Brenner:2024bps] for a specific example. Thus, the above Lagrangian may even be a complete EFT (to that order). However, it is best understood as a toy model, in which we can explore which types of effects may arise due to non-minimal interactions and we choose the model guided by technical simplicity, in particular requiring second-order equations of motion. As we will see, the model is already complex enough to exhibit several distinct effects in black-hole images.
@@
Previously, Ref. [Prasanna:2003ix] derived bounds on the non-minimal coupling $\alpha$. From the birefringent bending of light by the sun, Ref. [Prasanna:2003ix] obtained $\sqrt{|\alpha|}\lesssim 3.16\times 10^{5}\,\mathrm{km}$, while from radar echo experiments of radio signals passing in the vicinity of the sun they obtained $\sqrt{|\alpha|}\lesssim 6.25\times 10^{4}\,\mathrm{km}$, and $\sqrt{|\alpha|}\lesssim 2.45\,\mathrm{km}$ from the timing of the binary pulsar PSR B1534+12 signals. Taking these constraints at face value would rule out $|\alpha|/M^{2}\sim\mathcal{O}(1)$ (for $M$ the mass of a supermassive black hole), which is the range of values for which we expect black-hole images to change appreciably. These bounds should in our view be interpreted with a grain of salt for the following reason. Astrophysical constraints on couplings are typically subject to a lack of control over the system (in contrast to laboratory experiments), and the present constraints are no exception. For instance, in some astrophysical environments, both magnetic fields as well as a non-vanishing Ricci scalar are present. Both affect the propagation of light, if higher-order field-strength terms as well as a non-minimal Ricci coupling are included in the Lagrangian. Translating observations into constraints is subject to assumptions about whether or not such couplings are present in Lagrangians. Existing constraints from astrophysical systems are derived under simplifying assumptions, such as the absence of a standard Euler-Heisenberg term (higher-order field strength term), which modifies the propagation of light in the presence of magnetic fields.
However, our main motivation to consider $\alpha$ is not that we are interested in observational constraints on this specific coupling. Rather, we consider $\alpha$ as one representative of a much larger class of non-minimal light-curvature couplings, for which it has never been explored before what their effects on photon rings are. It is, however, crucial to understand their effects, because, if future observational data on photon rings is used to constrain new physics in gravity, it is not sufficient to only consider modifications of the spacetime geometry. Accounting for new-physics effects that change the propagation of light is also crucial. Here, we take a first step in this direction. In doing so, we focus on $\alpha$ specifically, rather than other (so far unconstrained) non-minimal couplings, purely for reasons of calculational convenience, because $\alpha$, being part of the Horndeski family and having an effective metric description for photon propagation, c.f. Sec. III, allows us to compute effects easily. We expect that the effects induced by $\alpha$ are paradigmatic of those created by non-minimal couplings more generally, motivating our study of their imprints on black-hole images.
@@ III Propagation of non-minimally coupled photons
In the presence of a non-minimal coupling, the propagation of photons is, in general, not described by the standard geodesic equation [Drummond:1979pp]. The latter is obtained from the wave-like solutions of motion for minimally coupled electromagnetic fields in the geometric-optics limit. The equation of motion is modified by non-minimal interactions; therefore, the null geodesics of the background spacetime $g_{\mu\nu}$ lose their interpretation as describing the propagation of photons.
At a technical level, some settings with non-minimal couplings enable a description through the geodesic equation of an effective metric $\tilde{g}_{\mu\nu}$. This effective metric does not encode the geometry of spacetime; instead, it captures the effect of the non-minimal interaction at an effective level. This can be done if there is a field transformation for the particular solution $g_{\mu\nu}\rightarrow\tilde{g}_{\mu\nu}$ that maps the Lagrangian in Eq. (1) into the Einstein-Maxwell Lagrangian plus higher-curvature corrections. However, as we will see below, the situation for black-hole backgrounds is more complex and, in particular, a single effective metric does not suffice.
@@
where the square brackets correspond to the anti-symmetrization of indices as $T_{[\mu\nu]}=(T_{\mu\nu}-T_{\nu\mu})/2$. Eq. (7) is a homogeneous system of equations for the tensor $f_{\mu\nu}$ that admits non-trivial solutions if and only if $k^{\mu}$ is suitably constrained. To extract the form of the relevant constraints, we note that the Riemann tensor for the Schwarzschild spacetime can be written as [Drummond:1979pp]
@@
where $\{U_{(i)}^{\mu\nu}\}_{i=1}^{2}$ are defined in terms of vierbein fields $\{e^{\mu}_{(a)}\}_{a=0}^{3}$ satisfying $g^{\mu\nu}=\eta^{ab}e^{\mu}_{(a)}e^{\mu}_{(b)}$, with $\eta^{ab}$ the inverse flat metric, as [Drummond:1979pp]
@@
It holds that $k^{\mu}l_{\mu}=0=k^{\mu}m_{\mu}$, so that $l_{\mu}$ and $m_{\mu}$ can be understood as the two real basis vectors which encode the linear polarization of a photon in the plane orthogonal to its direction of propagation. These orthogonal linear polarizations are chosen based on theoretical considerations, as they encapsulate the effects of the effective metrics induced by the non-minimal light-curvature coupling. However, linearly polarized radiation, as measured, for M87*, by the EHT Collaboration [EventHorizonTelescope:2021bee], is usually expressed as a combination of the three real Stokes parameters $\mathcal{I,Q,U}$ encoding the total intensity of the radiation and its linearly polarized components, see [Thompson:2017int]. While a mapping exists between the two descriptions, we will solely focus on the theoretical one provided by $l_{\mu}$ and $m_{\mu}$ in Eq. (11), as we will not compare our results with actual observational data on polarization. The corresponding effective metrics can be written (up to an arbitrary conformal factor, because the effective metrics only encode modified propagation of light) as [Chen:2015cpa]
@@
$1/2\leq\alpha/M^{2}\leq 1,$ (15)
in order for the effective metrics to be non-singular at the horizon $r=2M$ for both polarizations. Note that the radial coordinates of the metrics in Eqs. (4) and (12) have different physical interpretations. While in Eq. (4), $r$ corresponds to the geometrically meaningful areal radius, this is not the case for Eq. (12), whose areal radius is given by $r_{\rm areal}=r\sqrt{\rho^{(\bm{v})}(r)}$. Consequently, although the horizon is located at $r=2M$ in both cases, the areal radius of the horizon for photons in the effective metrics is $r_{\rm areal,H}=2M\sqrt{\rho^{(\bm{v})}(2M)}$.
In practice, the effective metrics above indicate that light rays do not propagate along null geodesics of the metric $g_{\mu\nu}$. Photons with different polarizations propagating on the Schwarzschild spacetime experience different effective metrics in the geometric optics approximation (later studies also considered the cases of charged and rotating black-hole spacetimes [Daniels:1993yi, Daniels:1995yw] and obtained the effective metrics in the geometric optics approximation, see, e.g., [Chen:2016hil]). This is a phenomenon known as bi-metricity [Visser:2002sf]. In the sections below, we will determine the impact this feature can have on black-hole images.